# Color code decoder with improved scaling for correcting circuit-level noise

### v1

@@ Abstract
Two-dimensional color codes are a promising candidate for fault-tolerant quantum computing, as they have high encoding rates, transversal implementation of logical Clifford gates, and high feasibility of magic state constructions. However, decoding color codes presents a significant challenge due to their structure, where elementary errors violate three checks instead of just two (a key feature in surface code decoding), and the complexity in extracting syndrome is greater. We introduce an efficient color-code decoder that tackles these issues by combining two matching decoders for each color, generalized to handle circuit-level noise by employing detector error models. We provide comprehensive analyses of the decoder, covering its threshold and sub-threshold scaling both for bit-flip noise with ideal measurements and for circuit-level noise. Our simulations reveal that this decoding strategy nearly reaches the best possible scaling of logical failure ($p_{\mathrm{fail}}\sim p^{d/2}$) for both noise models, where $p$ is the noise strength, in the regime of interest for fault-tolerant quantum computing. While its noise thresholds are comparable with other matching-based decoders for color codes ($8.2\%$ for bit-flip noise and $0.46\%$ for circuit-level noise), the scaling of logical failure rates below threshold significantly outperforms the best matching-based decoders.
@@ 1 Introduction
Two-dimensional (2D) color codes [1, 2] are a family of stabilizer quantum error-correcting codes that can be realized with local interactions on a 2D plane, and provide a promicing pathway for implementing fault-tolerant quantum computation. Compared to surface codes [3, 4], color codes have several noteworthy advantages: (i) They have higher encoding rates for the same code distance [5], (ii) all the Clifford gates can be implemented transversally [1], and (iii) an arbitrary pair of commuting logical Pauli product operators can be measured in parallel via lattice surgery [6]. In addition, the well-studied Steane code is a small instance of a color code, and recently a number of fault-tolerant operations including transversal logic gates and magic state injection have been demonstrated using color codes [7, 8, 9, 10, 11].
@@
Several approaches to decode errors in color codes have been proposed. The most widely studied methods are the projection decoder and its variants [12, 13, 14, 15, 16], which apply minimum-weight perfect matching (MWPM) on two or three restricted lattices (which are specific sub-lattices of the dual lattice of the color code) and then deduce the final correction by a procedure called ‘lifting’. These approaches can achieve thresholds of around 8.7% for bit-flip noise [12] and around 0.47% for circuit-level noise [16]. However, they have a fundamental limitation that the logical failure rate $p_{\mathrm{fail}}$ scales like $p^{d/3}$ below threshold [14, 17, 16], not $p^{d/2}$, where $p$ is a physical noise strength and $d$ is the code distance. This is because there exist errors with weights $O(d/3)$ that are uncorrectable via the decoder (noting that, in principle, an optimal decoder can correct any error with weight up to $d/2$). Such a drawback may significantly hinder resource efficiency of quantum computing, necessitating a larger code distance to maintain the same logical failure rate, compared to a scenario using a decoder with the expected optimal scaling of $p^{d/2}$. Roughly speaking, it demands $9/4$ times as many qubits if other factors are the same.
The Möbius MWPM decoder [17] is another matching-based decoder processed by applying the MWPM algorithm on a manifold built by connecting the three restricted lattices, which has the topology of a Möbius strip. It achieves a higher threshold of 9.0% under bit-flip noise and, more importantly, a better scaling of $p_{\mathrm{fail}}\sim p^{3d/7}$. The decoder has subsequently been improved to accommodate circuit-level noise and general color-code lattices [18].
@@ 2.1 Color codes
Figure 1: Example of a triangular color code and its restricted lattice. (a) The triangular color code with code distance $d=7$ based on a hexagonal lattice with three boundaries of different colors is displayed. Each hexagonal face is associated with a pair of $Z$-type and $X$-type checks, exemplified by $S_{f_{\texttt{r}}}^{Z}$ on a red face and $S_{f_{\texttt{b}}}^{X}$ on a blue face, respectively. Logical operators $X_{L}$ and $Z_{L}$ can be supported on one of the three boundaries. Each edge is colored by the color of the faces connected by it, as exemplified by the red edge highlighted. (b) The red-restricted lattice of the code is illustrated, where vertices and edges are shown as circles and purple lines, respectively. Its vertices consist of blue faces (blue circles), green faces (green circles), and an additional boundary vertex $v_{\mathrm{bdry}}$ (purple circle). Each of its edges corresponds to a red edge of the original lattice, which are highlighted as thick red lines.
@@ 2.3 Noise models and settings
To evaluate the performance of our decoder, we simulate $T$ rounds of the logical idling gate acting on the triangular color code of code distance $d$, which is preceded by logical initialization (to the $+1$ eigenstate |$\ket{\overline{0}}$ of $\overline{Z}$) and followed by a measurement in the $\overline{Z}$ basis. The logical initialization and measurement are done by initializing all the data qubits to |$\ket{0}$ and measuring them in the $Z$ basis, respectively. From this scenario, we can identify whether the $\overline{Z}$ observable fails (i.e., a $\overline{X}$ or $\overline{Y}$ error occurs) during the idling gate. While it is sufficient for bit-flip noise models, the $\overline{X}$ observable also can fail under circuit-level noise models. To cover this, we simulate again by swapping the $X$- and $Z$-type parts from the cnot schedule (see Sec. 4.3 for more details). The logical failure rate $p_{\mathrm{fail}}$ is computed as $p_{\mathrm{fail}}=p_{\mathrm{fail}}^{(X)}+p_{\mathrm{fail}}^{(Z)}-p_{\mathrm{fail}}^{(X)}p_{\mathrm{fail}}^{(Z)}$, where $p_{\mathrm{fail}}^{(X)}$ and $p_{\mathrm{fail}}^{(Z)}$ are respectively the failure rates of the $\overline{X}$ and $\overline{Z}$ observables, by assuming that these two observables fail independently.
@@ 2.4 Assessment of decoder performance
We assess the performance of the decoder from two perspectives: (i) what the noise threshold $p^{\times}_{\mathrm{}}$ of the decoder is and (ii) how the logical failure rate behaves when the noise strength $p$ is well below the threshold (namely, $p\lessapprox p^{\times}_{\mathrm{}}/2$).
@@
To investigate the sub-threshold scaling, we fit data into the ansatz
@@
Since $p_{\mathrm{fail}}/T$ is expected to be constant when $p=p^{\star}_{\mathrm{}}$, we regard $p^{\star}_{\mathrm{}}$ as a ‘noise threshold’ as well. To distinguish the two types of thresholds $p^{\times}_{\mathrm{}}$ and $p^{\star}_{\mathrm{}}$, we refer to them as the cross threshold and scaling threshold, respectively. Additionally, we specify the noise model (bit-flip or circuit-level) on the subscripts of the thresholds such as $p^{\times}_{\mathrm{bitflip}}$, $p^{\times}_{\mathrm{circuit}}$, $p^{\star}_{\mathrm{bitflip}}$, and $p^{\star}_{\mathrm{circuit}}$.
@@ 3 Concatenated MWPM decoder without faulty measurements
To formally describe the above procedure, let us first define some notations. For a lattice $\mathcal{L}$, the sets of its vertices, edges, and faces are respectively denoted as $\Delta_{0}\quantity(\mathcal{L})$, $\Delta_{1}\quantity(\mathcal{L})$, and $\Delta_{2}\quantity(\mathcal{L})$. For each color $\texttt{c}\in\quantity{\texttt{r},\texttt{g},\texttt{b}}$, we denote the sets of c-colored edges and faces in $\mathcal{L}_{\mathrm{2D}}$ as $\Delta_{1}^{(\texttt{c})}\quantity(\mathcal{L}_{\mathrm{2D}})$ and $\Delta_{2}^{(\texttt{c})}\quantity(\mathcal{L}_{\mathrm{2D}})$. The set of faces with violated $Z$-type checks is denoted as $\sigma_{Z}\in\Delta_{2}\quantity(\mathcal{L})$. We partition $\sigma_{Z}$ into $\sigma_{Z}=\sigma_{Z}^{(\texttt{r})}\cup\sigma_{Z}^{(\texttt{g})}\cup\sigma_{Z}^{(\texttt{b})}$ such that $\sigma_{Z}^{(\texttt{c})}\subseteq\Delta_{2}^{(\texttt{c})}\quantity(\mathcal{L}_{\mathrm{2D}})$ for each $\texttt{c}\in\quantity{\texttt{r},\texttt{g},\texttt{b}}$. The restricted and monochromatic lattices are then formally defined as follows:
@@
Upon the above definitions, the decoding process can be formally described as follows:
@@
      Perform the MWPM algorithm to identify a set $\widetilde{E}_{\mathrm{pred}}^{(\texttt{c})}\coloneqq\operatorname{MWPM}\quantity(\sigma_{Z}^{(\texttt{c}_{1})}\cup\sigma_{Z}^{(\texttt{c}_{2})};\mathcal{L}_{\neg\texttt{c}}^{*},v_{\mathrm{bdry}})\subseteq\Delta_{1}\quantity(\mathcal{L}_{\neg\texttt{c}}^{*})$, where $\texttt{c}_{1}$ and $\texttt{c}_{2}$ are the other two colors besides c. Define $E_{\mathrm{pred}}^{(\texttt{c})}\coloneqq\epsilon_{\neg\texttt{r}}^{-1}\quantity(\widetilde{E}_{\mathrm{pred}}^{(\texttt{c})})\in\Delta_{1}^{(\texttt{c})}\quantity(\mathcal{L}_{\mathrm{2D}})$.
@@
      Perform the MWPM algorithm to identify a set $\widetilde{V}_{\mathrm{pred}}^{(\texttt{c})}\coloneqq\operatorname{MWPM}\quantity(\sigma_{Z}^{(\texttt{c})}\cup E_{\mathrm{pred}}^{(\texttt{c})};\mathcal{L}_{\texttt{c}}^{*},v_{\mathrm{bdry}})\subseteq\Delta_{1}\quantity(\mathcal{L}_{\texttt{c}}^{*})$ and define $V_{\mathrm{pred}}^{(\texttt{c})}\coloneqq\epsilon_{\texttt{r}}^{-1}\quantity(\widetilde{V}_{\mathrm{pred}}^{(\texttt{c})})\in\Delta_{0}\quantity(\mathcal{L}_{\mathrm{2D}})$
@@
In Fig. 13 of Appendix C, we numerically verify that our color-selecting strategy of executing the sub-decoding procedures for all the three colors and selecting the best one is indeed necessary to maximize the performance of the decoder. If only one or two colors are considered in this process, the failure rate of the decoder significantly increases.
@@ 3.1 Performance analysis
We note that the value of $\beta$ is closely related to the least-weight uncorrectable errors; namely, a decoder may be able to correct any error of weight smaller than $O(\beta d)$ for a color code with code distance $d$. The projection decoder cannot correct specific types of errors with weights $O(d/3)$ [14, 17], which are correctable via the concatenated MWPM decoder as shown in Appendix B.1. Likewise, we can find weight-$O(3d/7)$ uncorrectable errors for the Möbius decoder [17].
One may be led to guess that the concatenated MWPM decoder can correct any error of weight smaller than $O(d/2)$. Surprisingly, however, there exist uncorrectable errors with weights $O(3d/7)$ just as with the Möbius decoder, as illustrated in Appendix B.2. Then how could we obtain $\beta\approx 1/2$ from the above numerical analysis? We speculate that small-weight uncorrectable errors with weights $<O(d/2)$ may exist only when $d$ is sufficiently large. As evidence, weight-$O(3d/7)$ uncorrectable errors of the type described in Appendix B.2 exist only when $d\geq 25$. Hence, we conjecture that $\beta$ may approach a limit of $3/7$ for sufficiently large values of $d$, although its confirmation would demand extensive computational resources. Nonetheless, the concatenated MWPM decoder is still beneficial compared to the Möbius decoder, where $\beta$ is computed to be about $3/7$ even for small values of $d<25$. Note that Ref. [17] also suggests a modification of the Möbius decoder that manually compares inequivalent recovery operators, which is proved to correct every error with weight less than $d/2$ when $d\leq 13$. This improvement has not been considered in Fig. 4.
An additional numerical analysis for bit-flip noise is presented in Fig. 13 of Appendix C, showing that our strategy of comparing the outcomes from the sub-decoding procedures of all the three colors is indeed effective compared to performing only one or two sub-decoding procedures.

### v2

@@ Abstract
Two-dimensional color codes are a promising candidate for fault-tolerant quantum computing, as they have high encoding rates, transversal implementation of logical Clifford gates, and resource-efficient magic state preparation schemes. However, decoding color codes presents a significant challenge due to their structure, where elementary errors violate three checks instead of just two (a key feature in surface code decoding), and the complexity of extracting syndrome is greater. We introduce an efficient color-code decoder that tackles these issues by combining two matching decoders for each color, generalized to handle circuit-level noise by employing detector error models. We provide comprehensive analyses of the decoder, covering its threshold and sub-threshold scaling both for bit-flip noise with ideal measurements and for circuit-level noise. Our simulations reveal that this decoding strategy nearly reaches the best possible scaling of logical failure ($p_{\mathrm{fail}}\sim p^{d/2}$) for both noise models, where $p$ is the noise strength, in the regime of interest for fault-tolerant quantum computing. While its noise thresholds are comparable with other matching-based decoders for color codes ($8.2\%$ for bit-flip noise and $0.46\%$ for circuit-level noise), the scaling of logical failure rates below threshold significantly outperforms the best matching-based decoders.
@@ 1 Introduction
Two-dimensional (2D) color codes [1, 2] are a family of stabilizer quantum error-correcting codes that can be realized with local interactions on a 2D plane, and provide a promising pathway to implementing fault-tolerant quantum computation. Compared to surface codes [3, 4], color codes have several noteworthy advantages: (i) They have higher encoding rates for the same code distance [5], (ii) all the Clifford gates can be implemented transversally [1], and (iii) an arbitrary pair of commuting logical Pauli product operators can be measured in parallel via lattice surgery [6]. Notably, the transversality of Clifford gates enables highly resource-efficient magic state preparation schemes without distillation [7, 8, 9], which can be concatenated with distillation to further improve the quality of the output magic state [10, 11].
Recent experiments have demonstrated significant advancements in the implementation of color codes. The distance-3 Steane code, a well-studied small instance of a color code, has been used to perform various fault-tolerant operations (such as transversal logic gates, magic state preparation, and lattice surgery) on trapped-ion [12, 13, 14], neutral atom [15, 16], and superconducting [17] platforms. Furthermore, Refs. [16, 17] also implement the distance-5 color code, with Ref. [17] demonstrating its improved logical error suppression compared to the Steane code.
@@
Several approaches to decode errors in color codes have been proposed. The most widely studied methods are the projection decoder and its variants [18, 19, 20, 21, 22], which apply minimum-weight perfect matching (MWPM) on two or three restricted lattices (which are specific sub-lattices of the dual lattice of the color code) and then deduce the final correction by a procedure called ‘lifting’. These approaches can achieve thresholds of around 8.7% for bit-flip noise [18] and around 0.47% for circuit-level noise [22]. However, they have a fundamental limitation that the logical failure rate $p_{\mathrm{fail}}$ scales like $p^{d/3}$ below threshold [20, 23, 22], not $p^{d/2}$, where $p$ is a physical noise strength and $d$ is the code distance. This is because there exist errors with weights $O(d/3)$ that are uncorrectable via the decoder (noting that, in principle, an optimal decoder can correct any error with weight up to $d/2$). Such a drawback may significantly hinder resource efficiency of quantum computing, necessitating a larger code distance to maintain the same $p_{\mathrm{fail}}$, compared to a scenario using a decoder with the expected optimal scaling of $p^{d/2}$. Roughly estimating based on the ansatz $p_{\mathrm{fail}}=\alpha(p/p_{\mathrm{th}})^{\beta d+\eta}$ for the threshold $p_{\mathrm{th}}$ and parameters $\alpha$, $\beta$, and $\eta$, using a decoder with $\beta=1/3$ demands about $9/4$ times more qubits than an optimal decoder with $\beta=1/2$ to achieve a similar $p_{\mathrm{fail}}$, assuming $p_{\mathrm{th}}$, $\alpha$, and $\eta$ remain similar in both cases.
The Möbius MWPM decoder [23] is another matching-based decoder implemented by applying the MWPM algorithm on a manifold built by connecting the three restricted lattices, which has the topology of a Möbius strip. It achieves a higher threshold of 9.0% under bit-flip noise and, more importantly, a better scaling of $p_{\mathrm{fail}}\sim p^{3d/7}$. The decoder has subsequently been improved to accommodate circuit-level noise and general color-code lattices [24].
@@ 2.1 Color codes
Figure 1: Example of a triangular color code and its restricted lattice. (a) The triangular color code with code distance $d=7$ based on a hexagonal lattice with three boundaries of different colors is displayed. Each face is associated with a pair of $Z$-type and $X$-type checks, exemplified by $S_{f_{\texttt{r}}}^{Z}$ on a red face and $S_{f_{\texttt{b}}}^{X}$ on a blue face, respectively. Logical operators $X_{L}$ and $Z_{L}$ can be supported on one of the three boundaries. Each edge is colored by the color of the faces connected by it, as exemplified by the red edge highlighted. (b) The red-restricted lattice of the code is illustrated, where vertices and edges are shown as circles and purple lines, respectively. Its vertices consist of blue faces (blue circles), green faces (green circles), and an additional boundary vertex $v_{\mathrm{bdry}}$ (purple circle). Each of its edges corresponds to a red edge of the original lattice, which are highlighted as thick red lines.
@@ 2.3 Noise models and settings
To evaluate the performance of our decoder, we simulate $T$ rounds of the logical idling gate acting on the triangular color code of code distance $d$, which is preceded by logical initialization (to the $+1$ eigenstate |$\ket{\overline{0}}$ of $\overline{Z}$) and followed by a measurement in the $\overline{Z}$ basis. The logical initialization and measurement are done by initializing all the data qubits to |$\ket{0}$ and measuring them in the $Z$ basis, respectively. From this scenario, we can identify whether the $\overline{Z}$ observable fails (i.e., a $\overline{X}$ or $\overline{Y}$ error occurs) during the idling gate. While it is sufficient for bit-flip noise models, the $\overline{X}$ observable also can fail under circuit-level noise models. To cover this, we simulate again by swapping the $X$- and $Z$-type parts from the cnot schedule (see Sec. 4.3 for more details). We estimate the logical failure rate $p_{\mathrm{fail}}$ as $p_{\mathrm{fail}}=p_{\mathrm{fail}}^{(X)}+p_{\mathrm{fail}}^{(Z)}$, where $p_{\mathrm{fail}}^{(X)}$ and $p_{\mathrm{fail}}^{(Z)}$ are respectively the failure rates of the $\overline{X}$ and $\overline{Z}$ observables. Since a $\overline{Y}$ error causes both observables to fail, this provides a conservative estimate that upper-bounds the actual failure rate.
@@ 2.4 Assessment of decoder performance
We assess the performance of the decoder from two perspectives: (i) what the noise threshold $p^{\times}_{\mathrm{}}$ of the decoder is and (ii) how the logical failure rate behaves when the noise strength $p$ is sufficiently lower than the threshold (namely, $p\lessapprox p^{\times}_{\mathrm{}}/2$).
@@
To investigate the sub-threshold scaling, we fit data (with sufficiently low $p$) into the ansatz222 This ansatz, which is widely employed in the literature [39, 40, 41, 42, 43, 23], can be justified as follows: For sufficiently low $p$, $p_{\mathrm{fail}}$ can be approximated by leaving only its lowest-order term with respect to $p$ (with no constant term, since $p_{\mathrm{fail}}=0$ when $p=0$). Furthermore, below the threshold, $p_{\mathrm{fail}}$ is expected to decrease exponentially with $d$, as proven for toric and surface codes [4, 44, 45, 46] and supported numerically for color codes [20, 23]. Hence, for sufficiently low $p$, it is natural to consider an ansatz where $\log p_{\mathrm{fail}}$ is bilinear in $\log p$ and $d$, which leads to the form of Eq. (3).
@@
Since $p_{\mathrm{fail}}/T$ is expected to be constant when $p=p^{\star}_{\mathrm{}}$, we regard $p^{\star}_{\mathrm{}}$ as a ‘noise threshold’ as well. To distinguish the two types of thresholds $p^{\times}_{\mathrm{}}$ and $p^{\star}_{\mathrm{}}$, we refer to them as the cross threshold and scaling threshold, respectively. Note that $p^{\times}_{\mathrm{}}$ and $p^{\star}_{\mathrm{}}$ are generally not equal, as the ansatz in Eq. (3) is an asymptotical expression valid for sufficiently small $p$, where the sub-leading order terms with respect to $p$ are negligible. Additionally, we specify the noise model (bit-flip or circuit-level) on the subscripts of the thresholds such as $p^{\times}_{\mathrm{bitflip}}$, $p^{\times}_{\mathrm{circuit}}$, $p^{\star}_{\mathrm{bitflip}}$, and $p^{\star}_{\mathrm{circuit}}$.
@@ 3 Concatenated MWPM decoder without faulty measurements
To formally describe the above procedure, let us first define some notations. For a lattice $\mathcal{L}$, the sets of its vertices, edges, and faces are respectively denoted as $\Delta_{0}\quantity(\mathcal{L})$, $\Delta_{1}\quantity(\mathcal{L})$, and $\Delta_{2}\quantity(\mathcal{L})$. For each color $\texttt{c}\in\quantity{\texttt{r},\texttt{g},\texttt{b}}$, we denote the sets of c-colored edges and faces in $\mathcal{L}_{\mathrm{2D}}$ as $\Delta_{1}^{(\texttt{c})}\quantity(\mathcal{L}_{\mathrm{2D}})$ and $\Delta_{2}^{(\texttt{c})}\quantity(\mathcal{L}_{\mathrm{2D}})$. The set of faces with violated $Z$-type checks is denoted as $\sigma_{Z}\in\Delta_{2}\quantity(\mathcal{L}_{\mathrm{2D}})$. We partition $\sigma_{Z}$ into $\sigma_{Z}=\sigma_{Z}^{(\texttt{r})}\cup\sigma_{Z}^{(\texttt{g})}\cup\sigma_{Z}^{(\texttt{b})}$ such that $\sigma_{Z}^{(\texttt{c})}\subseteq\Delta_{2}^{(\texttt{c})}\quantity(\mathcal{L}_{\mathrm{2D}})$ for each $\texttt{c}\in\quantity{\texttt{r},\texttt{g},\texttt{b}}$. The restricted and monochromatic lattices are then formally defined as follows:
@@
Using the above definitions, we formally describe the decoding process as follows:
@@
      Perform the MWPM algorithm to identify a set $\widetilde{E}_{\mathrm{pred}}^{(\texttt{c})}\coloneqq\operatorname{MWPM}\quantity(\sigma_{Z}^{(\texttt{c}_{1})}\cup\sigma_{Z}^{(\texttt{c}_{2})};\mathcal{L}_{\neg\texttt{c}}^{*},v_{\mathrm{bdry}})\subseteq\Delta_{1}\quantity(\mathcal{L}_{\neg\texttt{c}}^{*})$, where $\texttt{c}_{1}$ and $\texttt{c}_{2}$ are the other two colors besides c. Define $E_{\mathrm{pred}}^{(\texttt{c})}\coloneqq\epsilon_{\neg\texttt{c}}^{-1}\quantity(\widetilde{E}_{\mathrm{pred}}^{(\texttt{c})})\in\Delta_{1}^{(\texttt{c})}\quantity(\mathcal{L}_{\mathrm{2D}})$.
@@
      Perform the MWPM algorithm to identify a set $\widetilde{V}_{\mathrm{pred}}^{(\texttt{c})}\coloneqq\operatorname{MWPM}\quantity(\sigma_{Z}^{(\texttt{c})}\cup E_{\mathrm{pred}}^{(\texttt{c})};\mathcal{L}_{\texttt{c}}^{*},v_{\mathrm{bdry}})\subseteq\Delta_{1}\quantity(\mathcal{L}_{\texttt{c}}^{*})$ and define $V_{\mathrm{pred}}^{(\texttt{c})}\coloneqq\epsilon_{\texttt{c}}^{-1}\quantity(\widetilde{V}_{\mathrm{pred}}^{(\texttt{c})})\in\Delta_{0}\quantity(\mathcal{L}_{\mathrm{2D}})$
@@
In Fig. 14 of Appendix D, we numerically verify that our color-selecting strategy of executing the sub-decoding procedures for all the three colors and selecting the best one is indeed necessary to maximize the performance of the decoder. If only one or two colors are considered in this process, the failure rate of the decoder significantly increases.
@@ 3.1 Performance analysis
We note that the value of $\beta$ is closely related to the least-weight uncorrectable errors; namely, a decoder may be able to correct any error of weight smaller than $O(\beta d)$ for a color code with code distance $d$. The projection decoder cannot correct specific types of errors with weights $O(d/3)$ [20, 23], which are correctable via the concatenated MWPM decoder as shown in Appendix C.1. Likewise, we can find weight-$O(3d/7)$ uncorrectable errors for the Möbius decoder [23].
One may be led to guess that the concatenated MWPM decoder can correct any error of weight smaller than $O(d/2)$. Surprisingly, however, there exist uncorrectable errors with weights $O(3d/7)$ just as with the Möbius decoder, as illustrated in Appendix C.2. This raises the question of how the numerical analysis appears to scale as $\beta\approx 1/2$. We speculate that small-weight uncorrectable errors with weights $<O(d/2)$ may exist only when $d$ is sufficiently large. As evidence, weight-$O(3d/7)$ uncorrectable errors of the type described in Appendix C.2 exist only when $d\geq 25$. Hence, we conjecture that $\beta$ may approach a limit of $3/7$ for sufficiently large values of $d$, although its confirmation would demand extensive computational resources. Nonetheless, the concatenated MWPM decoder is still beneficial compared to the Möbius decoder, where $\beta$ is computed to be about $3/7$ even for small values of $d<25$. Note that Ref. [23] also suggests a modification of the Möbius decoder that manually compares inequivalent recovery operators, which is proved to correct every error with weight less than $d/2$ when $d\leq 13$. This improvement has not been considered in Fig. 4.
An additional numerical analysis for bit-flip noise is presented in Fig. 14 of Appendix D, showing that our strategy of comparing the outcomes from the sub-decoding procedures of all the three colors is indeed effective compared to performing only one or two sub-decoding procedures.